A laptop manufacturer wants to compare the total cost of assembling its laptops in the United States versus Taiwan. All of the laptops will be sold in the United States. To evaluate inventory, it uses a safety factor of 2.25. The holding cost per laptop is $4 per week in the United States and $3.50 per week in Taiwan. The lead time with U.S. production is one week, whereas it is eight weeks with production in Taiwan. In addition, it costs $2 to ship laptops to the United States from Taiwan. Weekly demand is 1,000 laptops, with a standard deviation of 800.

(Round your answer to 2 decimal places.)
a. What is the per unit holding cost of a laptop with U.S. production? per unit
(Round your answer to 2 decimal places.)
b. What is the per unit holding cost of a laptop with production in Taiwan? per unit
(Round your answer to 2 decimal places.)
c. A laptop requires 40 minutes of labor to assemble in the United States or in Taiwan. The total cost of labor in the United States is $40 per hour, while in Taiwan it is $10. What is the change in cost per laptop if it switches production to Taiwan? per unit
Answer #1

(a) The inventory on order with U.S. production is 1 * 1,000 = 1,000 units.

The inventory on-hand with U.S. production is square root of (1 + 1) * 1,000 * 2.25 = 3181.5 units.

Therefore, total average inventory is 1,000 + 3181.5 = 4181.5 units.

The average weekly holding cost is $4 * 4181.5 = $16726

Hence the average holding cost per unit with U.S. production is 16726 / 1,000 = $16.726.

Answer: $16.72.

(b)

The inventory on-order with that has production in Taiwan is 8 * 1,000 =8,000 units.

The inventory on-hand with production in Taiwan is square root of (8 + 1) *1,000 * 2.25 = 6,750 units.

Hence, the total average inventory is 8,000 + 6,750 =14,750.

Therefore, the average weekly holding cost is $3.50 * 14,750 = $51625.

The average holding cost per unit with production in Taiwan is 51625 / 1,000 = $51.62

Answer: $51.62

(c)

If production is in the United States, the labor cost per unit is $40 * 40/60 = $26.67.

However, there is no transportation cost, and the average holding cost per unit is $16.72. Therefore, the total cost is 26.67 + 16.726 = $43.39 per unit.

Now, if the production moves to Taiwan, the labor cost per unit is $10 * 40/60 = $6.7. However, unlike US, here there is a $2 per unit transportation cost and an average holding cost per unit of $51.62. This is a total cost of 6.7 + 2 + 51.62= $60.32 per unit.

The difference in the total costs is 60.32 - 43.39 = $16.93 units.

Therefore, the cost increases by $16.93 per unit if the production is switched to Taiwan.